The supply and demand equations for strawberries are given by: Qs= -10+5P Qd= 20-5P
a.What is the equilibrium market price and quantity for strawberries in this market?
b.Calculate the price elasticity of demand and price elasticity of supply at the equilibrium price.
1
Expert's answer
2016-12-19T09:01:11-0500
a) Qd=Qs, where Qs= -10+5P, Qd= 20-5P;
-10+5P=20-5P;
10P=30, P=3, Q=-10+5*3=5
Answer: P=3, Q=5.
b) Es = 5*(3/5)=3, ED=-(-5)*(3/5)=3.
Answer:Es=3, ED=3.